# Pathways to crystal chirality An algorithm to identify new displacive chiral phase transitions

### v1

@@ I Introduction
Chirality in periodic solids is characterized by structures belonging to space groups that include only operations of the first kind [1, 2, 3], i.e., those which map a right (left)-handed coordinate system to another right (left)-handed system. This fundamental property distinguishes chiral structures from their achiral counterparts [4, 5]. The concept of chirality in crystallography extends beyond mere point group symmetry and encompasses the complete three-dimensional periodicity of the crystal lattice [6]. Chiral space groups, also known as Sohncke groups, comprise 65 of the 230 crystallographic space groups and are characterized by the absence of improper symmetry operations such as inversion centers, mirror planes, and rotoinversion axes [4, 7, 1]. This inherent asymmetry gives rise to unique physical properties in chiral materials, including optical activity [8, 9] or nonlinear optical effects [10], which have significant implications in various fields ranging from materials science to pharmaceutical research [11]. Recently, the study of chiral effects in crystals has gained considerable attention, particularly concerning phenomena such as chiral phonons [12, 3], which are phonons that exhibit rotational motion (clockwise or counterclockwise) that are locked with the particle’s momentum. Another area of interest is spin selectivity [13], where one spin state is favored over another when electrons pass through chiral magnetic structures. Additionally, a recent research focus is polar and magnetic skyrmions [14], topologically protected textures of spin or electric dipoles. There is a growing interest in exploring topological states, such as those associated with Kramers-Weyl fermions [15], alongside the emerging field of altermagnetism [16, 17]. Chiral crystal structures, characterized by the absence of inversion symmetry and mirror planes, are particularly promising in this context. These unique structures can give rise to distinctive magnetic properties, positioning them as natural candidates for altermagnetic phenomena [16] while also profoundly influencing the electronic structure and band dispersion [18].
@@
In this study, we propose a systematic methodology for identifying achiral parent structures from known chiral crystalline phases belonging to the 22 enantiomorphic space groups and capable of exhibiting displacive phase transitions. We employ both rigorous symmetry analysis and first-principles calculations. For simplicity and clarity, we restrict our focus to non-magnetic crystals displaying chiral enantiomorphic space groups as magnetic chirality can be at play as well [3]. We use group-theoretical approaches to derive the parent structure from the atomic positions of the low-symmetry chiral phase. This methodology is designed for use in high-throughput computational searches, potentially uncovering a diverse range of new candidates for chiral transitions. Once the parent achiral structure has been identified relying on symmetry arguments, subsequent density functional theory (DFT) calculations are conducted to compute phonon instabilities and check the existence of soft modes connecting the achiral reference and the chiral structures. By presenting this framework, we aim to stimulate the scientific community to identify and investigate novel materials exhibiting chiral transitions. This could potentially lead to discoveries with implications for applications such as optoelectronics, spintronics, and nonlinear optics.
@@ II.2 The case of TeO2
In this section, we apply our methodology to the interesting case of TeO2 crystal. Among the three commonly reported polymorphs at ambient conditions [51, 52], this material has an enantiomorphic low-symmetry phase of either $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$ space groups. To our knowledge, no parent structure has yet been reported for this compound at the bulk level. Still, a rutile $P4_{2}/mnm$ high symmetry phase has been reported recently when grown epitaxially on FeTe substrate [53]. Our pseudosymmetry search predicts that the high symmetry $P4_{2}/mnm$ achiral structure can be linked to the chiral low-symmetry $P4_{1}2_{1}2$ (space group number 92) and $P4_{3}2_{1}2$ (space group number 96) phases through the $Z3$ irrep from the zone boundary $Z$ point, similarly to the case of K3NiO2[24]. Hence, we can check from DFT calculations whether a soft phonon mode is present at the $Z$ point of the $P4_{2}/mnm$ achiral phase of TeO2.
@@
The calculated structural data of the converged $P4_{2}/mnm$ structure computed with a $k$-mesh of 10$\times$10$\times$10 can be found in Tab. 1. We found one unstable mode of degeneracy two at the $Z$ point with a frequency of $370i$ cm-1. The symmetry of this unstable mode corresponds to the irrep $Z_{4}$, which is the right one to link the high-symmetry rutile phase and the low-symmetry chiral phases. Due to the double degeneracy of the $Z_{4}$ mode, different linear combinations $(a,b)$ can be constructed. Diagonal distortions $(a,a)/(-a,-a)$ and $(a,-a)/(-a,a)$ domains are associated to the $P4_{1}2_{1}2$ and $P4_{3}2_{1}2$ space groups respectively. Similarly to the case of K3NiO2, non-diagonal distortions $(a,b)$ drives the system to a $C222_{1}$ phase whenever $a\neq 0$ and $b\neq 0$ whereas $Cmcm$ phases are encountered when $a=0$ or $b=0$. The condensation and relaxation of the $Z_{4}$ unstable mode eigenvector give an energy gain of $\Delta E=-0.76$ eV, and $\Delta E=-0.33$ eV per 6 atoms unit cell for the $(a,a)$, $(a,0)$ phases respectively. We obtained that the $(a,b)$ case has the same energy gain as the $(a,a)$ case, so we can conclude that the case $a\neq b$ does not give a local minimum. The $(a,0)$ case has higher energy than the $(a,a)$ case such that we can conclude that, through the $Z_{4}$ irrep, the chiral $(a,a)$ domains are those giving the lowest energy.
@@
Table 1: Structural data in reduced coordinates for the high symmetry parent phase of TeO2 found with the pseudosymmetry search algorithm and after DFT relaxation. Unit cell parameters for the tetragonal structure are $a=b=5.060$Å, $c=3.401$Å. Forces and stresses where converged up to $10^{-6}$ Har/Bohr ($5.14\cdot 10^{-5}$ eV/Å). The last three columns represent the unrelaxed eigenvector displacements following the $Z_{4}$ mode towards space group number 92 ($P4_{1}2_{1}2$) in the home unit cell.
@@
A SAM analysis of our relaxed chiral phase reveals that the distortion for the $P4_{2}/mnm$ phase decomposes into $Z_{4}$ and $\Gamma_{1}$ irreps where the amplitude of the $\Gamma_{1}$ mode (0.004 Åper ion) is two orders of magnitude smaller than that of the $Z_{4}$ mode (0.093 Åper ion). The $\Gamma_{1}$ mode is a secondary SAM that appears through invariant coupling with the $Z_{4}$ unstable mode as this mode is not unstable at the zone center.
The $Z_{4}$ mode induces in-plane displacements of the Te atoms and both out-of-plane and in-plane displacements of the O atoms, although the latter is of lower magnitude. Similarly to K3NiO2, those distortions produce a helical structure as shown in Fig. 2. In Fig. 2, we show a schematic representation of the TeO2 crystal and its chiral $Z_{4}$ $(a,a)$ distortion. Atomic positions correspond to the high-symmetry $P4_{2}/mnm$ phase, and the arrows indicate how the atoms move when transitioning from the achiral rutile phase to the chiral $P4_{1}2_{1}2$ phase. The handedness of the chiral phase has been computed following the helicity method described in Ref. [54, 55]. This helicity is well-defined within the context of displacive phase transitions, allowing for a reference-independent comparison between different compounds. By normalizing its value to the number of atoms in the unit cell, the reference becomes intrinsically determined by the physical nature of the transition. For TeO2 a value of $7.38\cdot 10^{-3}$ is found, which is similar to the one found in K3NiO2 ($7.75\cdot 10^{-3}$) for a similar type of transition.
Figure 2: Schematic representation of the TeO2 crystal structure where atoms occupy the $P4_{2}/mnm$ high symmetry positions. Arrows indicate the direction of the atom displacements that bring the system to the $P4_{1}2_{1}2$ chiral phase. Gold and red balls represent Te and O atoms, respectively.
II.3 The case of Sr2As2O7
Here, we shall present another example of chiral phase transition predicted with the pseudosymmetry search algorithm. The Sr2As2O7 compound exhibits experimentally a low symmetry phase of either $P4_{1}$ or $P4_{3}$ chiral space groups [56, 57]. We could identify an achiral parent structure of $P4_{2}$ symmetry by applying the pseudosymmetry search algorithm. We have relaxed this new parent structure with a $k$-mesh of 4$\times$4$\times$2 and report the obtained structural parameters in Tab. 2.

### v2

@@ I Introduction
Chirality in periodic solids is characterized by structures belonging to space groups that include only operations of the first kind [1, 2, 3, 4, 5], i.e., those which map a right (left)-handed coordinate system to another right (left)-handed system. This fundamental property distinguishes chiral structures from their achiral counterparts [6, 7]. The concept of chirality in crystallography extends beyond mere point group symmetry and encompasses the complete three-dimensional periodicity of the crystal lattice [8]. Chiral space groups, also known as Sohncke groups, comprise 65 of the 230 crystallographic space groups and are characterized by the absence of improper symmetry operations such as inversion centers, mirror planes, and rotoinversion axes [6, 9, 1]. This inherent asymmetry gives rise to unique physical properties in chiral materials, including optical activity [10, 11] or nonlinear optical effects [12], which have significant implications in various fields ranging from materials science to pharmaceutical research [13]. Recently, the study of chiral effects in crystals has gained considerable attention, particularly concerning phenomena such as chiral phonons [14, 15, 16, 17], which are phonons that exhibit rotational motion (clockwise or counterclockwise) that are locked with the particle’s momentum. Another area of interest is spin selectivity [18], where one spin state is favored over another when electrons pass through chiral magnetic structures. Additionally, a recent research focus is polar and magnetic skyrmions [19], topologically protected textures of spin or electric dipoles. There is a growing interest in exploring topological states, such as those associated with Kramers-Weyl fermions [20], alongside the emerging field of altermagnetism [21, 22]. Chiral crystal structures, characterized by the absence of inversion symmetry and mirror planes, are particularly promising in this context. These unique structures can give rise to distinctive magnetic properties, positioning them as natural candidates for altermagnetic phenomena [21] while also profoundly influencing the electronic structure and band dispersion [23].
@@
In this study, we propose a systematic methodology for identifying achiral parent structures from known chiral crystalline phases belonging to the 22 enantiomorphic space groups and capable of exhibiting displacive phase transitions. We employ both rigorous symmetry analysis and first-principles calculations. For simplicity and clarity, we restrict our focus to non-magnetic crystals displaying chiral enantiomorphic space groups as magnetic chirality can be at play as well [3]. We use group-theoretical approaches to derive the parent structure from the atomic positions of the low-symmetry chiral phase. While illustrated here on a proof-of-concept set of six compounds, the methodology is well suited for future high-throughput implementation, offering the potential to uncover novel candidates for chiral transitions. Once the parent achiral structure has been identified relying on symmetry arguments, subsequent density functional theory (DFT) calculations are conducted to compute phonon instabilities and check the existence of soft modes connecting the achiral reference and the chiral structures. By presenting this framework, we aim to stimulate the scientific community to identify and investigate novel materials exhibiting chiral transitions. This could potentially lead to discoveries with implications for applications such as optoelectronics, spintronics, and nonlinear optics.
@@ II.2 The case of TeO2
In this section, we apply our methodology to the interesting case of TeO2 crystal. Among the three commonly reported polymorphs at ambient conditions [56, 57], this material has an enantiomorphic low-symmetry phase of either $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$ space groups. To our knowledge, no parent structure has yet been reported for this compound at the bulk level. Still, a rutile $P4_{2}/mnm$ high symmetry phase has been reported recently when grown epitaxially on FeTe substrate [58].
Figure 2: (a) Schematic representation of the TeO2 crystal structure where atoms occupy the $P4_{2}/mnm$ high symmetry positions. Arrows indicate the direction of the atom displacements that bring the system to the $P4_{1}2_{1}2$ chiral phase. Gold and red balls represent Te and O atoms, respectively. (b) Calculated phonon band structure along the $\Gamma-Z$ path of the high symmetry phase of TeO2. The color of the different branches are weighted according to the contribution of each chemical species to the dynamical matrix eigenvector (red for the O atom, and blue for the Te atom). (c) Energy surface when we condense the $Z_{3}$ unstable mode along the $(a,a)$ directions in normalized units.
Our pseudosymmetry search predicts that the high symmetry $P4_{2}/mnm$ achiral structure can be linked to the chiral low-symmetry $P4_{1}2_{1}2$ (space group number 92) and $P4_{3}2_{1}2$ (space group number 96) phases through the $Z3$ irrep from the zone boundary $Z$ point, similarly to the case of K3NiO2[29]. The calculated structural data of the converged $P4_{2}/mnm$ structure computed with a $k$-mesh of 10$\times$10$\times$10 can be found in the appendix.
@@
We can therefore study from DFT calculations whether a soft phonon mode is present at the $Z$ point of the $P4_{2}/mnm$ achiral phase of TeO2. As shown in Fig. 2(b), a double degenerate unstable phonon mode is encountered at $Z$, showing a frequency of $370i$ cm-1. The symmetry of this unstable mode corresponds to the irrep $Z_{3}$, which is the right one to link the high-symmetry rutile phase and the low-symmetry chiral phases. Due to the double degeneracy of the $Z_{3}$ mode, different linear combinations $(a,b)$ can be constructed. Diagonal distortions $(a,a)/(-a,-a)$ and $(a,-a)/(-a,a)$ domains are associated to the $P4_{1}2_{1}2$ and $P4_{3}2_{1}2$ space groups respectively. Similarly to the case of K3NiO2, non-diagonal distortions $(a,b)$ drives the system to a $C222_{1}$ phase whenever $a\neq 0$ and $b\neq 0$ whereas $Cmcm$ phases are encountered when $a=0$ or $b=0$. The condensation of the $Z_{3}$ unstable mode eigenvector give an energy gain of $\Delta E=-0.127$ eV per atom for the $(a,a)$ phase as shown in Fig. 2(c). We also obtained that the $(a,b)$ case has the same energy gain as the $(a,a)$ case, so we can conclude that the case $a\neq b$ does not give a local minimum. The $(a,0)$ case shows a higher energy ($\Delta E=0.06$ eV) than the $(a,a)$ case such that we can conclude that, through the $Z_{3}$ irrep, the chiral $(a,a)$ domains are those giving the lowest energy.
A SAM analysis of our relaxed chiral phase reveals that the distortion for the $P4_{2}/mnm$ phase decomposes into $Z_{3}$ and $\Gamma_{1}$ irreps where the amplitude of the $\Gamma_{1}$ mode (0.004 Åper ion) is two orders of magnitude smaller than that of the $Z_{3}$ mode (0.093 Åper ion). The $\Gamma_{1}$ mode is a secondary SAM that appears through invariant coupling with the $Z_{3}$ unstable mode as this mode is not unstable at the zone center.
The $Z_{3}$ mode induces a slight in-plane displacements of the Te atoms, as well as both out-of-plane and in-plane displacements of the O atoms. However, the in-plane component of the O displacement is smaller than the out-of-plane one. As shown in the weighted phonon band structure of Fig. 2(b) and the magnitude of the arrows in Fig. 2(a) the distortion is mostly dominated by the O movements. Similarly to K3NiO2, the associated distortions produce a helical structure as shown in Fig. 2(a). In Fig. 2(a), we show a schematic representation of the TeO2 crystal and its chiral $Z_{3}$ $(a,a)$ distortion. Atomic positions correspond to the high-symmetry $P4_{2}/mnm$ phase, and the arrows indicate how the atoms move when transitioning from the achiral rutile phase to the chiral $P4_{1}2_{1}2$ phase. The handedness of the chiral phase has been computed following the helicity method described in Ref. [59, 60]. This helicity is well-defined within the context of displacive phase transitions, allowing for a reference-independent comparison between different compounds. By normalizing its value to the number of atoms in the unit cell, the reference becomes intrinsically determined by the physical nature of the transition. For TeO2 a value of $7.38\cdot 10^{-3}$ is found, which is similar to the one found in K3NiO2 ($7.75\cdot 10^{-3}$) for a similar type of transition.
II.3 The case of Sr2As2O7
Here, we shall present another example of chiral phase transition predicted with the pseudosymmetry search algorithm. The Sr2As2O7 compound exhibits experimentally a low symmetry phase of either $P4_{1}$ or $P4_{3}$ chiral space groups [61, 62]. We could identify an achiral parent structure of $P4_{2}$ symmetry by applying the pseudosymmetry search algorithm. We have relaxed this new parent structure with a $k$-mesh of 4$\times$4$\times$2 and the obtained structural parameters can be found in the appendix.
As in the previous case, a symmetry analysis of the supergroup–subgroup relationship reveals that this parent structure can be connected to the enantiomorphic space groups $P4_{1}$ and $P4_{3}$ through two different irreps at the $Z$ point. Indeed a SAM analysis gives $Z_{3}$ (with an amplitude of 0.054 Å per ion) and $\Gamma_{1}$ (with an amplitude of 0.016 Å per ion) for the $P4_{1}$ space group and $Z_{4}$ and $\Gamma_{1}$ for the $P4_{3}$ space group but with the same SAM amplitudes as $Z_{3}$ and $\Gamma_{1}$ in the $P4_{1}$ case. As for TeO2, the $\Gamma_{1}$ mode appears to be secondary while the $Z_{3}$ and $Z_{4}$ are the primary main modes. Unlike K3NiO2 and TeO2, where the transition to chiral enantiomorphic phases is driven by the same irreducible representation (irrep), the two chiral space groups in this case are associated with two distinct irreps. Our phonon calculations in the $P4_{2}$ achiral reference of Sr2As2O7 are consistent with these results as two unstable modes are found, one with an imaginary frequency of $-503i$ cm-1 and the $Z_{3}$ label and the other one with an imaginary frequency of $-501i$ cm-1 and the $Z_{4}$ label as shown in Fig. 4. These two modes of different symmetry have the same frequency, and condensing their respective eigenvectors and relaxing the structure gives the same gain of energy ($\Delta E=-0.18$ eV per atom), as we would expect for two pairs of enantiomorphic crystals. The substantial unstable frequencies and significant energy gain suggest that achieving experimental stability for the achiral parent structure may be challenging. However, these findings could motivate further experimental efforts to enhance the stability of this phase, potentially through the application of tensile strain or by optimizing growth conditions, as was successfully demonstrated for TeO2 [58].
This new example is interesting because it shows that our methodology can find displacive chiral phase transitions in which two different phonon modes are responsible for the two enantiomorphic space groups.
Figure 4: (a) Schematic representation of the Sr2As2O7 crystal structure where atoms occupy the $P4_{2}$ high symmetry positions. Arrows indicate the direction of the atom displacements that bring the system to the $P4_{3}$ chiral phase. Green, yellow, and red balls represent Sr, As, and O atoms, respectively. (b) Calculated phonon band structure along the $\Gamma-Z$ path of the high symmetry phase of Sr2As2O7. The color of the different branches are weighted according to the contribution of each chemical species to the dynamical matrix eigenvector (red for the O atom, green for the Sr and blue for the As atoms). (c) Energy surface when we condense the $Z_{4}$ unstable mode in normalized units.
In the limit of our knowledge, such a possibility has not been previously discussed in the literature.
II.4 The case of Na2SeO9
We shall now discuss the case of Na2SeO9. This material is only reported in the Materials Project and with the $P4_{1}22$ and $P4_{3}22$ pair of enantiomorphic space groups [63]. While no experimental report exists for this crystal, DFT calculations predict it to be stable. Given the scarcity of reported crystals within the enantiomorphic space groups, we can use it as a new test case for our search method for displacive chiral phase transition. After applying the pseudosymmetry search and relaxing the structures with a $k$-mesh sampling of 4$\times$4$\times$4, we obtain a high symmetry parent structure with the space group $P4_{2}/mmc$ and with structural data reported in the appendix, that can be connected to the enantiomorphic space groups by phonon modes at the $Z$ point.
By performing the phonon calculations at the $Z$-point, we found several unstable modes, as shown in Fig. 5. The irreps of the two most unstable modes are $Z_{1}$ and $Z_{2}$, which gives a nonchiral phase transition to space group $Pmmm$ and $Pccm$ respectively, with a degenerate frequency of $-237$ cm-1. The third and fourth modes correspond to the $Z_{3}$ and $Z_{4}$ irreps, which are also degenerate and drive a chiral phase transition showing the same frequency of $-217$ cm-1 and similar symmetry. In contrast to TeO2 or K3NiO2, diagonal distortion domains $(a,a)$ or $(-a,-a)$ in this system lead to a $Pmma$ achiral space group symmetry, whereas distortions along $(a,0)$ or $(0,a)$ drive the chiral transition towards the $P4_{1}22$ and $P4_{3}22$ space groups, respectively.
Figure 5: (a) Schematic representation of the Na2SeO9 crystal structure where atoms occupy the high symmetry positions ($P4_{2}/mmc$ space group). Arrows indicate the direction of the atom displacements that bring the system to the $P4_{3}22$ chiral phase. Gold, green, and red balls represent Na, Se, and O atoms. (b) Calculated phonon band structure along the $\Gamma-Z$ path of the high symmetry phase of Na2SeO9. The color of the different branches are weighted according to the contribution of each chemical species to the dynamical matrix eigenvector (red for the O atom, green for the Na and blue for the Se atoms). (c) Energy surface when we condense the $Z_{3}$ unstable mode in normalized units.
Upon relaxing the previous unstable modes, we observe that, despite exhibiting the highest instability, the $Z_{1}$ and $Z_{2}$ modes do not lead to the most stable phase, as they provide an energy gain of $\Delta E_{1}=-0.007$ eV per atom versus $\Delta E_{2}=-0.45$ eV per atom for the $Z_{3}$ and $Z_{4}$ modes in the chiral phase. A SAM decomposition of the relaxed $Z_{3}$ mode in $(0,a)$ domain gives $\Gamma_{1}$ (with amplitude 0.15Åper ion) and $Z_{3}$ (with amplitude 0.30Åper ion). The relaxation of the $Z_{3}$ and $Z_{4}$ modes along the diagonal domains $(a,a)$ or $(-a,-a)$ with the achiral $Pmma$ phase gives an energy gain of $E_{3}=-0.18$ eV per atom, which is larger than the relaxation coming from the $Z_{1}$ and $Z_{2}$ modes but lower than the chiral phase. Hence, Na2SeO9 has the potential to exhibit a displacive chiral phase transition.
II.5 Other materials showing no displacive chiral transitions
Figure 6: Calculated phonon band structure along the $\Gamma-Z$ path of the high symmetry phase of As2O5. The color of the different branches are weighted according to the contribution of each chemical species to the dynamical matrix eigenvector (red for the O atom, and blue for the As atoms).
First, let us consider the case of As2O5. The reported low symmetry chiral phase corresponds to a $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$ space group [64], and we found a parent phase of symmetry $P4_{2}/mnm$ following a pseudosymmetry search approach. From symmetry considerations, we found that the irrep that can connect with the chiral phase should be observed at the $Z$ point.
Table 1: Summary table collecting chiral space groups, achiral parent structures found with pseudosymmetry, irreps, and corresponding domains driving the phase transition, frequency of the instabilities, energy gains and helicity. The  ✗  sign indicates that either no parent phase was found or no soft phonon modes driving the transition were identified. Energy differences between high-symmetry and chiral phases after relaxation of the eigendisplacements are given per atom.
Compound Chiral group Achiral Ref. Irrep $\nu$ (cm-1) E (eV/atom) $\mathcal{H}$
TeO2 $P4_{1}2_{1}2$ $P4_{3}2_{1}2$ $P4_{2}/mnm$ $Z_{4}~{}(a,a)$ $Z_{4}~{}(a,-a)$370 -0.13 $7.38\cdot 10^{-3}$
Sr2As2O7 $P4_{1}$ $P4_{3}$ $P4_{2}$ $Z_{3}$ $(a)$ $Z_{4}~{}(a)$503 -501 -0.18 $0.81\cdot 10^{-3}$
Na2SeO9 $P4_{1}22$ $P4_{3}22$ $P4_{2}/mmc$ $Z_{3}$ $(a,0)$ $Z_{3}~{}(0,a)$217 -0.45 $2.70\cdot 10^{-2}$
As2O5 $P4_{1}2_{1}2$ $P4_{3}2_{1}2$ $P4_{2}/mnm$ $Z_{4}~{}(a,a)$ $Z_{3}~{}(a,a)$
Rb2Be2O3 $P4_{1}2_{1}2$ $P4_{3}2_{1}2$
CaTe2O3 $P4_{1}2_{1}2$ $P4_{3}2_{1}2$
However, unlike previous cases, no unstable mode is found when we compute the phonons at the $Z$ point as it is shown in Fig. 6. This suggests that displacive chiral phase transitions in this system are not physically present, at least at the bulk level.
Another reason a material may not exhibit a displacive chiral phase transition is because it is impossible to identify any achiral parent structures through a pseudosymmetry search. We found that this is the case for the Rb2Be2O3 compound that presents $P4_{1}2_{1}2$ or $P4_{3}2_{1}2$ low-symmetry chiral phases [65] or CaTeO3 that shows $P4_{1}$ or $P4_{3}$ chiral space groups [66]. In this second situation, we have a clear symmetry demonstration that no achiral phase can be built from the chiral one.
III Discussion
In this work, we have proposed a method combining pseudosymmetry search and first-principles calculations to identify achiral parent structures from known chiral phases connected through a soft phonon mode as present in displacive phase transitions. Our methodology, schematized in Fig. 1, helps discover new candidates for such phase transitions and disregards those where such a transition is impossible. In Table 1, we summarize the different cases where we have applied our methodology collecting the most important results of the study. The methodology helps to identify whether a soft phonon mode with the right symmetry can connect the achiral and chiral phases through continuous and relatively small displacements, e.g., TeO2, Sr2As2O7 or Na2SeO9. It also helps to disregard the cases where the pseudosymmetry would predict a possible achiral-to-chiral connection according to the symmetry. Still, no unstable mode is observed in the achiral phase from the DFT calculation, so it is physically improbable that a displacive phase transition exists, e.g., in the case of As2O5. Although, other mechanisms not captured by the present workflow, such as electronic instabilities like the Jahn-Teller one, could also potentially lead to chiral phase transitions [31]. Finally, it also helps to quickly show that no achiral phase can be constructed from the pseudosymmetry approach, which guarantees the absence of any smooth achiral to chiral transition as no high symmetry phase can be constructed, e.g., the cases of CaTeO3 and Rb2Be2O3. We could identify three new possible candidates (TeO2, Sr2As2O7 or Na2SeO9) that can have a soft phonon mode connecting the achiral and the chiral phase. This suggests that many more materials may display this behavior beyond the few examples reported in the literature so far. Interestingly, Sr2As2O7 appears to belong to a new class of systems where different irreps govern the chiral transitions to each of the enantiomorphic space groups. This discovery underscores the potential of our algorithm to reveal novel and intriguing chiral transitions. Further theoretical investigations would be valuable to determine whether similar behavior exists in other compounds, and to explore the possibility of additional new mechanisms for chiral transitions, thus opening exciting avenues for future research.
Our methodology provides a systematic and accurate way to screen various materials to search for new candidates for displacive chiral phase transitions. As only one soft-mode driven chiral transition has been reported within the 22 enantiomorphic space groups, this methodology is well-suited to be implemented on a high-throughput scheme and check whether the lack of reported structures is due to an overlooked problem or if it is related to some chemical or physical reason [3]. We hope this work will inspire the community to explore new soft-mode-driven chiral phase transitions together with the non-linear phononics ones [16] and to motivate experimentalists to check our predictions.
IV Methods
IV.1 Computational details
V Data Availability
All data regarding the structural properties of the high symmetry and low symmetry phases are provided in the appendix and in the dedicated supplementary file in plain text format.
Acknowledgements.
Appendix A Structural data
In this appendix we report the structural data of the high-symmetry phases encountered for the TeO2, Na2SeO9 and Sr2As2O7 crystals together with their associated displacements towards their respective chiral phases as reported in the main text.
@@ Appendix A Structural data
Table 2: Structural data in reduced coordinates for the high symmetry parent phase of TeO2 found with the pseudosymmetry search algorithm and after DFT relaxation. Unit cell parameters for the tetragonal structure are $a=b=5.060$Å, $c=3.401$Å. Forces and stresses where converged up to $10^{-6}$ Har/Bohr ($5.14\cdot 10^{-5}$ eV/Å). The last three columns represent the unrelaxed eigenvector displacements following the $Z_{3}$ mode towards space group number 92 ($P4_{1}2_{1}2$) in the home unit cell.
@@
Atom Wyckoff x y z $\delta_{x}$ $\delta_{y}$ $\delta_{z}$
Na 4b 0.5000 0.5000 0.0000 0.0000 0.1447 0.0000
Na 4a 0.0000 0.5000 0.5000 0.0000 -0.0322 0.0000
Se 4b 0.5000 0.0000 0.0000 0.0000 0.0474 0.0000
O 4c 0.5000 0.5000 0.6250 0.0749 0.0749 0.0000
O 8d 0.0000 0.1753 0.2500 -0.1075 0.0597 0.0109
O 8d 0.0000 0.5000 0.1471 -0.0814 -0.0164 -0.0206
O 8d 0.2808 0.1448 $0$0.0416 -0.2538 -0.0589
O 8d 0.2808 0.1448 0.0000 0.1342 -0.1193 -0.0131